OpenStudy (anonymous):

A used motorcycle is on sale for $3,600. Erik makes an offer equal to 3/4 of this price. How much does Erik offer for the motorcycle?

OpenStudy (camerondoherty):

Multiply \[\frac{ 3,600 }{ 1 }\times \frac{ 3 }{ 4 }\]

OpenStudy (camerondoherty):

This was my 1000th question answered by the way! c:

OpenStudy (skullpatrol):

:O Congrats!

OpenStudy (camerondoherty):

Thank You! cx